Find the measures of two complementary angles if one angle is 14 degrees more than 3 times the other angle.
Answer by nerdybill(7384)

Find the measures of two complementary angles if one angle is 14 degrees more than 3 times the other angle
.
Let x = one angle
then
3x+14 = second angle
.
definition of complementary angles:
x + 3x+14 = 90
4x+14 = 90
4x = 76
x = 19 degrees (one angle)
.
second angle:
3x+14 = 3(19)+14 = 71 degrees